OpenStudy (anonymous):

A skier is gliding along at 2.0m/s on horizontal, frictionless snow. He suddenly starts down a 15∘ incline. His speed at the bottom is 13m/s what is the length of the incline? How long does it take him to reach the bottom?

OpenStudy (jfraser):

set up a table with the pieces of information: \(\Delta x\), \(v_i\), \(v_f\), a, and t

OpenStudy (anonymous):

F=ma F=m*g*sin(15) a=F/m=g*sin(15)=2.54 m/s^2 v^2=u^2+2as s=(v^2-u^2)/2a s=(13^2-2^2)/(2*2.54) s=32.53m